The sum of two numbers, x and y, is 12. The difference of x and two times y is 6. What are the values of x and y?

Answer by stanbon(75887)

The sum of two numbers, x and y, is 12. The difference of x and two times y is 6. What are the values of x and y?
-----------------
Equations:
x + y = 12
x - y = 6
------------
Add and solve for "x":
2x + 18
x = 9
----
y = 12-x = 3
--------------------
Cheers,
Stan H.
====================